Consider the following abbreviated financial statements for Barrie Enterprises:

 

BARRIE Enterprises
2017 and 2018 Partial Statement of Financial Position
AssetsLiabilities and Owner's Equity
  2017 2018  2017 2018
Current assets$922$998Current liabilities$369$401
Net fixed assets 3,847 4,616Long-term debt 2,031 2,161

 

BARRIE Enterprises
2018 Statement of Comprehensive Income
Sales$12,000
Costs 5,445
Depreciation 1,073
Interest paid 302

 

a. What is owner's equity for 2017 and 2018? (Omit $ sign in your response.)

 

  
Owner's equity 2017 Numeric ResponseEdit Unavailable. 2369 correct.
Owner's equity 2018 Numeric ResponseEdit Unavailable. 3052 correct.

 

b. What is the change in net working capital for 2018? (Omit $ sign in your response.)

 

Change in NWC           $  Numeric ResponseEdit Unavailable. 44 correct.

 

c1. In 2018, Barrie Enterprises purchased $1,934 in new fixed assets. How much in fixed assets did Barrie Enterprises sell? (Omit $ sign in your response.)

 

Fixed assets sold           $ 

 

c2. In 2018, Barrie Enterprises purchased $1,934 in new fixed assets. What is the cash flow from assets for the year? (The tax rate is 35%.) (Omit $ sign in your response.)

 

Cash flow from assets           $ 

 

d1. During 2018, Barrie Enterprises raised $386 in new long-term debt. How much long-term debt must Barrie Enterprises have paid off during the year? (Omit $ sign in your response.)

 

Debt retired           $ 

 

d2. During 2018, Barrie Enterprises raised $386 in new long-term debt. What is the cash flow to creditors? (Omit $ sign in your response.)

 

Cash flow to creditors           $ 
Answer #1

Owner's equity = Total Assets - Total liability

For 2017 -

Total asset = Current asset + Net fixed asset

= 948+4107

= $5055

Total liability = Current liability + long term debt

= 382 + 2161

= $2543

Owner's equity = 5055 - 2543

= 2512

For 2018 -

Total asset = Current asset + Net fixed asset

= 1024 + 4876

= $5900

Total liability = Current liability + long term debt

= 416+2304

= $2720

Owner's equity = 5900 - 2720

= $3180

Fixed asset sold in 2018 = Net asset in 2017 - net asset in 2017 - depreciation + purchase

= 4107 - 1203 - 4876 +2077

= $105


answered by: ANURANJAN SARSAM